OpenStudy (anonymous):

What polynomial has roots of -5,-2, and 4?

OpenStudy (anonymous):

(x+5)(x+2)(x-4)

OpenStudy (amistre64):

there are an infinite number of them

OpenStudy (anonymous):

(x+5)(x+2)(x-4 x^3+3 x^2-18 x-40

OpenStudy (anonymous):

Add any constant outside the parentheses to make another polynomial with those zeros.

OpenStudy (amistre64):

add a constant is prolly a poor choice of words tho :)

OpenStudy (anonymous):

amistre64, you are correct. A better wording is multiply the expression by any constant.

OpenStudy (anonymous):

Ishaan94 is correct. The reason why is because the root is x = some number. If we want x = -5, x = - 2, and x = 4, then we need to be able to rewrite this as an equation of form (x+a) = 0, where a is the root. So we have x = -5, which becomes x+5 = 0, x = -2 becomes x + 2 = 0, and x = 4 becomes x - 4 = 0. This is the reason why these are called the zeroes of x. Another name yet is the x-intercept, as it is where the graph of the function will cross the x-axis. Hope this helps. http://www.tutorsean.net

OpenStudy (anonymous):

You are geniuses

OpenStudy (across):

\[f(x)=x3+6x^2−18x+C\]Where C is any constant. Therefore, as amistre pointed out, there's an infinite number of solutions to this question. ;p